please help me analyze this spectroscopy question
MF: CH..0 H-NMR 13C-NMR 0 0 4 - . 1 0 5
Answer #1

. Cg H140 (DBE = Home 8 - 14+1 = 9-1=2 82.2Cq, 2a) A SISS (5,34 1 Homa 51.58 (85, 311) - sl.62 (s. 314/ 82.1(513H) 55.8 (tiza